# Content Aware Analysis of Scholarly Networks: A Case Study on CORD19 Dataset

### v1

@@ Abstract
This paper investigates the relationships among key elements of scientific research network, namely articles, researchers, and journals. We introduce a novel approach to use semantic information through the HITS algorithm based propagation of topic information in the network. The topic information is derived by using the Named Entity Recognition and Entity Linkage. In our case, MedCAT is used to extract the topics from the CORD19 Dataset, which is a corpus of academic articles about COVID-19 and coronavirus scientific network. Our approach focuses on the COVID-19 domain, utilizing the CORD-19 dataset to demonstrate the efficacy of integrating topic-related information within the citation framework. Through the application of a hybrid HITS algorithm, we show that incorporating topic data significantly influences article rankings, revealing deeper insights into the structure of the academic community.
@@ 1 Introduction
Gaining insight into the relationship between the core elements of scientific research can be useful for a variety of purposes, such as guiding scientific effort toward better use of resources, inferring comparative results between fields of research, better representing the importance of certain research fields and research groups.
@@
Current scientific literature continues to grow at a rapid pace every day. Let alone being able to follow the growth of communities in which we are not a part, it has become very difficult to even find conferences, journals or other prominent studies in our field. Naturally, examining the academic community in detail becomes a great burden for most young or experienced researchers, which results in missing out promising researchers and useful works. To overcome this problem, most researchers represent the scientific literature as a wide network consists of different entities such as researchers, institutes, etc. In this paper, we aimed to analyze current literature and demonstrate different approaches to this problem with some practical applications.
@@
The academic writers, their studies and the citation connection between them composes the scientific community, which forms a wide network of authors and articles. Authors are identified as the entities that creates the knowledge in the community through the articles they have published. The citation network which is derived from the published work is the most common representation of this knowledge, which is very simple yet effective to analyze the communities. Social Network Analysis is a way of measuring and mapping various aspects of relationships between different entities such as people, organizations and groups Sweet [2018]. At first step, we started our analysis from simple representation of the network which is a graph of authors and articles. Then, we focus on the possible interpretations of the centrality metrics, PageRank and its variations in the real world scenarios.
@@
Apart from the approach above, it is clear that the proposed citation network lacks of the semantic meaning of the published works. Also representation of the topics is missing in the constructed graph of authors and their articles. Even though some solutions based on Natural Language Processing are available in the literature, most of these works require to process and analyze the content of the published articles via their open access files or abstracts. So, we propose a new and robust method to represent and retrieve the data in scientific network by considering the topics as an entity in the research graph. The topics are derived from a pipeline based on Named Entity Recognition and Knowledge Base of the relevant graph. Naturally, we have to focus on a specific domain to use the knowledge base effectively. Eventually, we showed the applicability of our method on a chosen domain and dataset, which are COVID-19 and CORD19 Dataset.
@@
Our motivation in this paper is to develop a way to provide the researchers with quantifiable information about the relationships between these elements so that it can be used for such purposes. This quantifiable information includes graph measures of individual elements of a graph as well as the graph measures of the whole graph. Briefly, we propose a pipeline to create, analyze and store the research network which consists of authors, articles, named entities and relationships between them.
@@ 2 Related Works
The idea of Social Network Analysis was firstly proposed in 1969 by Philip Mayer and Julia C. Mitchell in their work on Urban situation in Central African Towns Mayer and Mitchell [1970]. Over time, Social Network Analysis has been used to understand relationships between different groups, organizations, communities, and any other possible network Su et al. [2019]. Social Network analysis has a interdisciplinary nature as being in the intersection of sociology, mathematics, statistics and computer science. In addition to these, the rise of Big Data in recent years lead to analyze communities in more efficient way. For example, it has been a common approach to analyze connected social media data to detect misinformation or influential behaviours Ahmed et al. [2020], Vassey et al. [2022].
Scientific community is another domain which can be analyzed to derive meaningful information because of its interconnected nature. One of the earliest work in the literature is the Impact Factor proposed by Garfield (1972) in order to calculate the effectiveness and impact of the journals Garfield [1972]. He calculated the ImpactFactor with the formula:
@@
where $A$ is the number of times articles published in journal j in years $i-1$ and $i-2$ were cited in indexed journals and $B$ is the number of articles, reviews, proceedings or notes published in journal j years $i-1$ and $i-2$. Following his works many researcher aimed to rank articles, authors and journals based on their impact on the scientific community Cai et al. [2018].
Considering the article ranking methods, plenty of earlier works are derived from PageRank algorithm Page et al. [1999]. Mostly this approach cause biased results due to the fact that the older papers ,which have naturally higher citations than the newer ones, are assigned with higher ranking. CiteRank was proposed to remove the bias caused by PageRank through assigning more probability to new articles so that random surfer model can choose these articles Walker et al. [2006]. Additionally, FutureRank Sayyadi and Getoor [2009] and P-Rank Yan et al. [2011] algorithms were proposed in order to make use of different aspects such as time-indicator, authorship, journal information, etc. P-Rank hifted the focus to the understanding heterogeneous network representation of the entities Zhang et al. [2019]. The heterogeneous network proposed in the P-Rank algorithm consists of author, article and journal layers which propagates information among themselves to rank the specific article in the main network.
Apart from these works, HITS algorithm was proposed by Kleinberg. HITS algorithm uses authority and hub concepts to exploit local structure of the network Kleinberg [1999]. The W-Rank then used the both PageRank and HITS algorithms in order to utilize link weights based on citation and authorship relationships Zhang et al. [2019].
The existing solutions in the literature generally ignore the importance of the different edges in the heterogeneous network. Some researchers use topics for academic search by using topic modelling and its integration into the random walk framework Tang et al. [2008], however most of these methods lack of motivation to use topics in a weighting scheme to understand the nature of the community. Even though the time information is also used to evaluate link importance in citation relations in CiteRank algorithm Walker et al. [2006], it suffers from the absence of semantic meaning and heterogeneity. Also we firstly present a way to asses the semantic meaning through approaching to the topics as not only a similarity measure but also a part of the network.
In this paper, we introduce the use of the MedCAT Concept Annotation Tool Kraljevic et al. [2020] and relevant knowledge base in the coronavirus domain. Based on the heterogeneous network built with academic entities, we have conducted bunch of experiments for the different link weighting schemes, whereas proposing a different approach for ground truth.
@@ 3 Methodology
  • Old articles tend to have higher citation, which leads to biased results in ranking algorithms.
@@
  • The prestigious authors tend to publish articles with the bigger influence.
  • Important articles are cited by other important articles. The meaning of all the citations are not the same.
  • The semantically similar papers tend to cite each others and such a citation has more importance if their topics are dominant in the network.
Considering these assumptions, our network structure mainly based on the citation network which is composed of article nodes. The authors of and journal, if possible, of the article are connected to it with author and journal vertices. These attributes can be though as another layer of the network which is used to propagate information in prior works. The time information is also used to weight citation link between articles. Such network structure is very similar to PageRank + HITS approaches we mentioned above with various weigthing schemes. Differently, we add the topic layer, similar to author and journal layers, to the network by also exploring the influence of topics on citation weights. Eventually, a lightweight semantic network can be a part of the graph, which helps to analyze semantic relationships between articles in the network. We also believe that such as approach enables us to exploit topic based search and understand topic-wise prestige of articles in the network.
Before starting to theoretical groundings of the algorithm, we are stating that the works Wang et al. [2013] and Zhang et al. [2019] are used as a basis for our PageRank + HITS algorithm. Our main contribution is introducing the concept of the topic for such an algorithm.
@@ 3.1 Heterogeneous Network
where $V_{tp}$ represents the vertices of the topic network, whereas $E_{ar-tp}$ stands for the edges from articles to topics. $E_{tp-tp}$ is a term added for topic network which has a hierarchy derived from the thesaurus and ontology of biomedical concepts tanks to Unified Medical Language System (UMLS).
@@ 3.2.1 Article-Author
It is quite obvious that authors contribute differently to their published works. Although in some disciplines such as computer science, the ordering of the authors implies the importance and contribution, it is not a standardized approach in scientific community. Additionally, it can lead to underestimating the contribution of the authors, which is a serious issue in academic community. Some research includes $H-Index$ based solutions however this approach can assign lower ranks young researchers which have lower $H-Index$.
@@ 3.2.2 Article-Journal
The journals tends to publish similar quality articles in line with their own prestige, so we believe that we reach the articles published with equal probability starting from a specific journal. There are only two possibility in the article-journal networks, which are publish or not publish Zhang et al. [2019].
@@ 3.2.3 Article-Article
The citation network has more information than other sub-networks, which stems from the complex and versatile nature of the citation relationship. In addition to the graph based approach, many works explore the Natural Language Processing based techniques to understand the importance of the citation. However most of these works require huge amount of effort and data to train and classify the relevant models to understand whether a citation is influential or not. We use a hybrid approach by combining graph attributes and processing the abstract information to find similarities between articles. Using the abstract is more robust, efficient and fast than trying to find citation text in the document. Also it prevents us from being limited to articles that are only open access PDF. Briefly, the citation weight should have two different parts which are semantic-based similarity and network-based similarity Zhang et al. [2019]. Based on the work proposed by Zhang et al. we improve the network-based similarity by adding parameters related to authorship and journals in which published.
@@ 3.3 Topic Linking and Semantic Weighting
Having a graph-structured representation of the research world allows the addition of explicit connections to other graph-structured knowledge representations. One prime example of such knowledge representations is ontologies. The possible connections between a selected ontology, Unified Medical Language System (UMLS), have been investigated. The connections between the papers and the UMLS concepts are constructed by passing the abstracts of the papers through a named entity recognizer called MedCAT.
In this paper, we propose a method that is based on Named Entity Recognition and Linking (NER+L) to extract the relevant concepts from the article abstracts based on MedCAT and Unified Medical Language System.
MedCAT Kraljevic et al. [2020] is an content annotation tool based on Word2Vec embeddings which can be used to extract information from medical documents to link them to medical ontologies such as UMLS Lindberg et al. [1993].
@@ 3.4 Ranking Algorithm
Based on the HITS algorithm Kleinberg [1999], we use a weighted iteration and updates of authorities and hubs in the network. The previous research Wang et al. [2013] and Zhang et al. [2019] are followed when defining the scores in the algorithm. Naturally, due to the new introduced topic network, the score functions are a slightly different.
@@ 3.4.1 Hub Scores
The hub scores in the scientific network can be interpreted as the quality and impact of the hubs which the relevant entity belongs to. We analyzed the hub scores of articles, authors, journals and topics in this direction. We followed the slight derivation of HITS algorithm by normalizing the hub score with the number of the links Wang et al. [2013], because of the risk that authors and journal which publish huge number of articles can dominates the hubs. In addition to work of Wang et al., we present the hub scores of the topics by considering the topics as a part of the heterogeneous network. Apart from these, to understand current state of the network we also followed the time-aware approaches similar to prior works Wang et al. [2013]. Time-aware weights for article-author, article-journal and article-article links are needed to score the hubs.
@@ 3.4.2 Authority Scores
The algorithm in the previous work Zhang et al. [2019] is employed to calculate the authority scores of articles with the score functions above.
@@ 4.1 Dataset
We use the public dataset CORD19 which is a corpus of academic articles about COVID-19 and coronavirus scientific network published by Semantic Scholar Team at the Allen Institute for AI Wang et al. [2020]. The final version was released on June 2, 2022 with 1M+ papers and around 370k full text support. We have performed a pre-processing that eliminates the duplicate articles, articles with no year information or journal information and broken rows. Also the dataset has not any information about the citations among the articles. To create a graph representation of CORD19 Dataset, we have used Semantic Scholar API to fetch the citation information related to each articles. By doing this operation, we also aimed to synchronize dataset with the Semantic Scholar Academic Graph API. Eventually, we have created and pıblished a graph dataset containing 728675 articles, 2210182 authors, and 5875663 citations. The dataset represents the citation network of the CORD19 Dataset on January 2024.
For experiments it would be computationally challenging to use all the graph to find the the related metrics, so we have performed our analysis on a subset of the graph containing 19981 articles, 121431 author, 2925 journals, and 209788 citations.
The semantic network was created using the MedCAT Kraljevic et al. [2020] outputs of the article abstracts. Here we follow the Unified Medical Language System Lindberg et al. [1993] and the unique identifiers in the Metathesaurus Corpus. The extracted Concept Unique Identifiers (CUI) are used to build the semantic network by creating the links to related article by setting a accuracy threshold, %75 in our pipeline. Some Concept Unique Identifiers are discarded by considering the their type ids, see Appendix C and D.
@@ 4.1.1 Ground Truth and Evaluation Criteria
Finding a ground truth to analyze the academic networks is a common issue that most researchers tried to answer Su et al. [2019]. In our paper, we do not follow the prior ground truth because it would be contradictory to use metrics and algorithms based on old approaches as ground truth while trying to find a novel metric or trying to show that new parameters change the results. Since there was no human-based annotation on the dataset we worked with, we made a comparative evaluation criterion. In this case, we tried to analyze how different parameters in the algorithm change the result and what different semantic network causes.

### v2

@@ Abstract
This paper investigates the relationships among key elements of the scientific research network, namely articles, researchers, and journals. We introduce a novel approach to use semantic information through the HITS algorithm-based propagation of topic information in the network. The topic information is derived by using the Named Entity Recognition and Entity Linkage. In our case, MedCAT is used to extract the topics from the CORD19 Dataset, which is a corpus of academic articles about COVID-19 and the coronavirus scientific network. Our approach focuses on the COVID-19 domain, utilizing the CORD-19 dataset to demonstrate the efficacy of integrating topic-related information within the citation framework. Through the application of a hybrid HITS algorithm, we show that incorporating topic data significantly influences article rankings, revealing deeper insights into the structure of the academic community.
@@ 1 Introduction
Gaining insight into the relationship between the core elements of scientific research can be useful for a variety of purposes, such as guiding scientific efforts toward better use of resources, inferring comparative results between fields of research, and better representing the importance of certain research fields and research groups.
@@
Current scientific literature continues to grow at a rapid pace every day. Let alone being able to follow the growth of communities of which we are not a part, it has become very difficult to even find conferences, journals, or other prominent studies in our field. Naturally, examining the academic community in detail becomes a great burden for most young or experienced researchers, which results in missing out on promising researchers and useful works. To overcome this problem, most researchers represent the scientific literature as a wide network consisting of different entities such as researchers, papers, and journals with the help of some appropriate network measures [3], [19], [21], [23]. In this paper, we aimed to analyze current literature and demonstrate different approaches to this problem with some practical applications.
@@
The academic writers, their studies, and the citation connection between them compose the scientific community, which forms a wide network of authors and articles. Authors are identified as the entities that create knowledge in the community through the articles they have published. The citation network which is derived from the published work is the most common representation of this knowledge, which is very simple yet effective in analyzing the communities. Social Network Analysis is a way of measuring and mapping various aspects of relationships between different entities such as people, organizations, and groups [15]. In the first step, we started our analysis from a simple representation of the network which is a graph of authors and articles. Then, we focus on the possible interpretations of the centrality metrics, PageRank, and its variations in real-world scenarios.
@@
Apart from the approach above, it is clear that the proposed citation network lacks the semantic meaning of the published works [11]. Also, the representation of the topics is missing in the constructed graph of authors and their articles. Even though some solutions based on Natural Language Processing are available in the literature, most of these works require processing and analyzing the content of the published articles via their open-access files or abstracts [11], [18]. So, we propose a new and robust method to represent and retrieve the data in the scientific network by considering the topics as an entity in the research graph. The topics are derived from a pipeline based on Named Entity Recognition and Knowledge Base of the relevant field, which requires a focus on a specific domain to use the knowledge base effectively. Eventually, we showed the applicability of our method on a chosen domain and dataset, which are the COVID-19 and CORD19 Dataset.
@@
Our motivation in this paper is to develop a way to provide the researchers with quantifiable information about the relationships between these elements so that they can be used for such purposes. This quantifiable information includes graph measures of individual elements of a graph as well as the graph measures of the whole graph. Briefly, we propose a pipeline to create, analyze, and store the research network which consists of authors, articles, named entities, and relationships between them.
Moreover, we introduce the use of the MedCAT Concept Annotation Tool [7] and relevant medical ontology thanks to UMLS [8]. Based on the heterogeneous network built with academic entities, we have conducted a bunch of experiments for the different link weighting schemes, on the other hand proposing a different approach for ground truth.
@@ 2 Related Works
The idea of Social Network Analysis was first proposed in 1969 by Philip Mayer and Julia C. Mitchell in their work on Urban situations in Central African Towns [9]. Over time, Social Network Analysis has been used to understand relationships between different groups, organizations, communities, and any other possible network [14]. Social Network analysis has an interdisciplinary nature as is in the intersection of sociology, mathematics, statistics, and computer science. In addition to these, the rise of Big Data in recent years has led to the analysis of communities in a more efficient way. For example, it has been a common approach to analyze connected social media data to detect misinformation or influential behaviors [1], [17].
The scientific community is another domain that can be analyzed to derive meaningful information because of its interconnected nature. One of the earliest works in the literature is the Impact Factor proposed by Garfield (1972) in order to calculate the effectiveness and impact of the journals [4]. He calculated the ImpactFactor with the formula:
@@
where $A$ is the number of times articles published in journal j in years $i-1$ and $i-2$ were cited in indexed journals and $B$ is the number of articles, reviews, proceedings or notes published in journal j years $i-1$ and $i-2$. Following his works many researchers aimed to rank articles, authors, and journals based on their impact on the scientific community [3].
Considering the article ranking methods, plenty of earlier works are derived from the PageRank algorithm [12]. Mostly this approach causes biased results due to the fact that the older papers, which have naturally higher citations than the newer ones, are assigned with higher ranking. CiteRank was proposed to remove the bias caused by PageRank by assigning more probability to new articles so that the random surfer model can choose these articles [19]. Additionally, FutureRank [13] and P-Rank [22] algorithms were proposed in order to make use of different aspects such as time indicator, authorship, journal information, etc. P-Rank shifted the focus to understanding heterogeneous network representation of the entities [23]. The heterogeneous network proposed in the P-Rank algorithm consists of author, article, and journal layers which propagate information among themselves to rank the specific article in the main network.
Apart from these works, the HITS algorithm was proposed by Kleinberg. HITS algorithm uses authority and hub concepts to exploit the local structure of the network [6]. The W-Rank then used both PageRank and HITS algorithms in order to utilize link weights based on citation and authorship relationships [23].
The existing solutions in the literature generally ignore the importance of the different edges in the heterogeneous network. Some researchers use topics for academic search by using topic modeling and its integration into the random walk framework [16], however most of these methods lack of motivation to use topics in a weighting scheme to understand the nature of the community. Even though the time information is also used to evaluate link importance in citation relations in CiteRank algorithm [19], it suffers from the absence of semantic meaning and heterogeneity. Also, we first present a way to assess the semantic meaning by approaching the topics as not only a similarity measure but also a part of the network.
@@ 3 Methodology
  • Old articles tend to have higher citations, which leads to biased results in ranking algorithms.
@@
  • The prestigious authors tend to publish articles with a bigger influence.
  • Important articles are cited by other important articles. The meaning of all the citations is not the same.
  • The semantically similar papers tend to cite each other and such a citation has more importance if their topics are dominant in the network.
Considering these assumptions, our network structure is mainly based on the citation network which is composed of article nodes. The authors of and journal, if possible, of the article are connected to it with author and journal vertices. These attributes can be thought of as another layer of the network that is used to propagate information in prior works. The time information is also used to weight citation links between articles. Such network structure is very similar to the PageRank + HITS approaches we mentioned above with various weighting schemes. Differently, we add the topic layer, similar to the author and journal layers, to the network by also exploring the influence of topics on citation weights. Eventually, a lightweight semantic network can be a part of the graph, which helps to analyze semantic relationships between articles in the network. We also believe that such an approach enables us to exploit topic-based search and understand the topic-wise prestige of articles in the network.
Before starting to theoretical groundings of the algorithm, we are stating that the works [21] and [23] are used as a basis for our PageRank + HITS algorithm. Our main contribution is introducing the concept of the topic for such an algorithm.
@@ 3.1 Heterogeneous Network
where $V_{tp}$ represents the vertices of the topic network, whereas $E_{ar-tp}$ stands for the edges from articles to topics. $E_{tp-tp}$ is a term added for the topic network which has a hierarchy derived from the thesaurus and ontology of biomedical concepts tanks to Unified Medical Language System (UMLS).
Figure 1: A schematic of the heterogeneous network
@@ 3.2.1 Article-Author
It is quite obvious that authors contribute differently to their published works. Although in some disciplines such as computer science, the ordering of the authors implies the importance and contribution, it is not a standardized approach in the scientific community. Additionally, it can lead to underestimating the contribution of the authors, which is a serious issue in the academic community. Some research includes $H-Index$ based solutions however this approach can assign lower ranks to young researchers who have lower $H-Index$.
@@ 3.2.2 Article-Journal
The journals tend to publish similar quality articles in line with their own prestige, so we believe that we reach the articles published with equal probability starting from a specific journal. There are only two possibilities in the article-journal networks, which are publish or not publish [23].
@@ 3.2.3 Article-Article
The citation network has more information than other sub-networks, which stems from the complex and versatile nature of the citation relationship. In addition to the graph-based approach, many works explore Natural Language Processing techniques to understand the importance of the citation. However, most of these works require a huge amount of effort and data to train and classify the relevant models to understand whether a citation is influential or not. We use a hybrid approach by combining graph attributes and processing the abstract information to find similarities between articles. Using the abstract is more robust, efficient, and fast than trying to find citation text in the document. Also, it prevents us from being limited to articles that are only open-access PDFs. Briefly, the citation weight should have two different parts which are semantic-based similarity and network-based similarity [23]. Based on the work proposed by Zhang et al. we improve the network-based similarity by adding parameters related to authorship and journals in which published.
@@ 3.3 Topic Linking and Semantic Weighting
Having a graph-structured representation of the research world allows the addition of explicit connections to other graph-structured knowledge representations. One prime example of such knowledge representations is ontologies. The possible connections between a selected ontology and, the Unified Medical Language System (UMLS), have been investigated. The connections between the papers and the UMLS concepts are constructed by passing the abstracts of the papers through a named entity recognizer called MedCAT.
In this paper, we propose a method that is based on Named Entity Recognition and Linking (NER+L) to extract the relevant concepts from the article abstracts based on MedCAT and the Unified Medical Language System.
MedCAT [7] is a content annotation tool based on Word2Vec embeddings that can be used to extract information from medical documents to link them to medical ontologies such as UMLS [8].
@@ 3.4 Ranking Algorithm
Based on the HITS algorithm [6], we use a weighted iteration and updates of authorities and hubs in the network. The previous research [21] and [23] are followed when defining the scores in the algorithm. Naturally, due to the newly introduced topic network, the score functions are slightly different.
@@ 3.4.1 Hub Scores
The hub scores in the scientific network can be interpreted as the quality and impact of the hubs to which the relevant entity belongs. We analyzed the hub scores of articles, authors, journals, and topics in this direction. We followed the slight derivation of the HITS algorithm by normalizing the hub score with the number of the links [21], because of the risk that authors and journals that publish huge numbers of articles can dominate the hubs. In addition to the work of Wang et al., we present the hub scores of the topics by considering the topics as a part of the heterogeneous network. Apart from these, to understand the current state of the network we also followed the time-aware approaches similar to prior works [21]. Time-aware weights for article-author, article-journal, and article-article links are needed to score the hubs.
@@ 3.4.2 Authority Scores
The algorithm in the previous work [23] is employed to calculate the authority scores of articles with the score functions above.
@@ 4.1 Dataset
We use the public dataset CORD19 which is a corpus of academic articles about COVID-19 and coronavirus scientific network published by Semantic Scholar Team at the Allen Institute for AI [20]. The final version was released on June 2, 2022 with 1M+ papers and around 370k full text support. We have performed a pre-processing that eliminates the duplicate articles, articles with no year information or journal information and broken rows. Also the dataset has not any information about the citations among the articles. To create a graph representation of CORD19 Dataset, we have used Semantic Scholar [5] to fetch the citation information related to each articles. By doing this operation, we also aimed to synchronize dataset with the Semantic Scholar Academic Graph API. Eventually, we have created and published a graph dataset containing 728675 articles, 2210182 authors, and 5875663 citations. The dataset represents the citation network of the CORD19 Dataset on January 2024.
For experiments, it would be computationally challenging to use all the graphs to find the the related metrics, so we have performed our analysis on a subset of the graph containing 19981 articles, 121431 authors, 2925 journals, and 209788 citations, by considering the papers with citation 20 or more. After paper selection, we investigate the citations among these new paper sets. This process can be though of as a cutting of papers with low citations. The resulting graph is not so dense because only links between them are considered.
The semantic network was created using the MedCAT [7] outputs of the article abstracts. Here we follow the Unified Medical Language System [8] and the unique identifiers in the Metathesaurus Corpus. The extracted Concept Unique Identifiers (CUI) are used to build the semantic network by creating the links to related articles by setting an accuracy threshold, %75 in our pipeline. Some Concept Unique Identifiers are discarded by considering their type ids, see Appendix C and D.
@@ 4.1.1 Ground Truth and Evaluation Criteria
Finding a ground truth to analyze the academic networks is a common issue that most researchers tried to answer [14]. In our paper, we do not follow the prior ground truth because it would be contradictory to use metrics and algorithms based on old approaches as ground truth while trying to find a novel metric or trying to show that new parameters change the results. Since there was no human-based annotation on the dataset we worked with, we made a comparative evaluation criterion. In this case, we tried to analyze how different parameters in the algorithm change the result and what different semantic network causes.
@@
We assess the correlation through Spearman’s rank correlation [10]:
$\rho=\frac{\sum_{i}\left(R_{1}(P_{i})-\overline{R}_{1}\right)\left(R_{2}(P_{i})-\overline{R}_{2}\right)}{\sqrt{\sum_{i}\left(R_{1}(P_{i})-\overline{R}_{1}\right)^{2}\sum_{i}\left(R_{2}(P_{i})-\overline{R}_{2}\right)^{2}}}$
where $R(P_{i})$ are position of a specific in each lists, $\overline{R}$ are average ranks of papers in these lists. When there is a tie, the rank position is the average rank of all the ties.
Table 1: The top 10 ranked articles
Topic Author Journal Article PageRank